OpenStudy (anonymous):

21 is what percent of 168? proportion

OpenStudy (anonymous):

21/168 x 100 = 12.5 Therefore, 21 is 12.5 percent, or one eighth, of 168.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

so .... should I stop replying in the other thread? are you interested in learning how to do this or do you just want a list of answers?

OpenStudy (anonymous):

no im listening

OpenStudy (anonymous):

okay so let's say you wanted to know what percent 5 was of 30. \[ {5 \over 30} = {x \over 100} \] do you understand that so far? 5 out of 30 is equal to somethig out of 100

OpenStudy (anonymous):

Ok so would you write 4% of 350 like this? 4 __ 360 = n _____ 100

OpenStudy (anonymous):

350*

OpenStudy (anonymous):

almost -- if you start out knowing the percentage, then you're asking a slightly different question -- "4 out of 100" (that's 4%) is equal to how many out of 350? \[ {4 \over 100} = {x \over 350} \]

OpenStudy (anonymous):

oh.. ty

OpenStudy (anonymous):

the percentage is always the one over 100 -- it might be the one you know, or the one you're solving for.

OpenStudy (anonymous):

4% of 350= 350(.o4)=14

OpenStudy (anonymous):

depending on the problem. For example, let's say someone asked 36 is 56% of what? That's like saying "56 out of 100, is equal to 36 out of something", so it would be \[ {36 \over x} = {56 \over 100} \]

OpenStudy (anonymous):

The proportion is always 4 numbers, but the one that's "x" just changes based on what piece of information is unknown, that you're trying to compute.

OpenStudy (anonymous):

is that making more sense now?

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

a percent is always out of 100, so make sure the percentage you're given is the one over 100 if you're given a percentage. Or, make sure that x is the thing over 100 if the percentage is the thing you're solving for.

OpenStudy (anonymous):

how would you write this one, 80% of what number is 15

OpenStudy (anonymous):

that's just like the example I gave you above. can you try using that?

OpenStudy (anonymous):

first, how would you write 80% as a percentage?

OpenStudy (anonymous):

I mean, as a proportion?

OpenStudy (anonymous):

um n/15=80/100?

OpenStudy (anonymous):

almost -- you got the 80/100 right, that's 80 percent. but 80% is equal to 15 out of some number, so it's 15/n.